# Grokking Group Multiplication with Cosets

### v1

@@ Acknowledgements
We would like to thank Coreweave for donating the computing resources that we used to run all of our experiments and to Bilal Chughtai for helpful discussions we had throughout the project. We would also like to thank Nora Belrose, Aidan Ewart, Hailey Schoelkopf, and Cédric Simal for their helpful feedback on an earlier draft of this paper.
@@ Appendix A Author Contributions
Wrote the code for training and for calculating the Group Fourier Transform over $S_{n}$. Performed the initial analyses of models trained on $S_{5}$ and initially found what we came to call the coset circuit. Designed and ran causal interchange interventions. Derived formal properties of the coset circuit. Participated in discussions throughout the project and in writing the paper.
@@
Appendix B Experiment Details
We conducted experiments focusing on the permutation group of $S_{5}$ and $S_{6}$.
Model Epoch Linear Layer Size Embedding Size Optimizer Learning Rate
small 250,000 128 256 Adam 0.001
medium 250,000 1024 512 Adam 0.001
large 250,000 4096 512 Adam 0.001
Table 3: $S_{5}$ model information
Model Epoch Linear Layer Size Embedding Size Optimizer Learning Rate
medium 200000 256 512 Adam 0.001
medium 200000 1024 512 Adam 0.001
large 200000 4096 512 Adam 0.001
Table 4: $S_{6}$ model information
Appendix C Extra Graphs
C.1 Correlation of the Unembedding Vectors
Figure 7: Correlation of the columns of an unembedding matrix, ordered so that columns that act on neurons of the same coset circuit type are next to each other. This illustrates both the amount of redundancy, as very few cosets only get a single neuron, and the degree to which the unembedding relies solely on constructive interference. The bulk of the action of the unembedding is just adding together the neurons from the same coset that have correlation near one. There are only two sets of neurons that exhibit destructive interference, which is identified by correlation near $-1$.
C.2 $(H_{1},H_{1})$ Circuit
(a) Left Permutation Pre-activations
(b) Right Permutation Pre-activations
(c) Activations of all permutations combinations (pre-ReLU)
(d) Activations of all permutations combinations (post-ReLU)
Figure 8: An example of a coset $(H_{1},H_{1})$ circuit.
C.3 Distribution over Subgroups and Cosets
Figure 9: Distribution of coset circuits (as a percentage of neurons in the linear layer) for 128 runs trained on $S_{5}$.
@@ D.3 Normal Subgroups
The double cosets of a normal subgroup is simply the usual cosets.
@@ D.5 An important case
D.6 Representation Theory
Definition D.17.
@@ D.6 Representation Theory
In other words, a representation maps a group element $g$ to a linear operator $f(g):V\rightarrow V$ where $V$ is a vector space of dimension $d$, so that the group multiplication becomes compositions of linear operators ($f(g\cdot h)=f(g)\circ f(h)$). Without explicit specifications, all representations in this paper are assumed to be over complex numbers.
@@
Example D.26.
The permutation representation is reducible, because there is a subspace of the $\mathbb{C}^{3}$ that is invariant to it’s action.
@@
As it turns out, there are no irreducible representations of $S_{3}$ that are three-dimensional. The largest irrep of $S_{3}$ is $\rho_{(2,1)}$, which is made of $2\!\times\!2$.
(2 1 3)213(2\;1\;3)( 2 1 3 ) $\mapsto\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}$
(3 2 1)321(3\;2\;1)( 3 2 1 ) $\mapsto\begin{pmatrix}\frac{1}{2}&-\frac{\sqrt{3}}{2}\\ \frac{\sqrt{3}}{2}&-\frac{1}{2}\end{pmatrix}$
(2 3 1)231(2\;3\;1)( 2 3 1 ) $\mapsto\begin{pmatrix}-\frac{1}{2}&\frac{\sqrt{3}}{2}\\ -\frac{\sqrt{3}}{2}&-\frac{1}{2}\end{pmatrix}$
@@
Appendix E Fourier transform over finite groups
@@
For more details and applications, one can refer to, for example, [15]. We would like to note that there is also an inverse transform that restores the original function $f$ from $\hat{f}$:
@@
E.1 Projection on Fourier Basis
The inputs to the model are two one-hot vectors, $x_{l},\;x_{r}$, which multiply the embedding matrices $E_{l}x_{l}$ and $E_{r}x_{r}$. $E_{l}$ and $E_{r}$ are $d\times|G|$ matrices, where $d$ is the embedding dimension and $|G|$ is the size of the group. The columns are the embedding vectors for a single element $g\in G$. The normal approach would be to try and look at the column spaces of $E_{l}$ and $E_{r}$, as these columns are the inputs to the model. However, since each row of $E_{l}$ and $E_{r}$ and each value of that row is associated with a single element of $G$, we instead treat each row of the embedding as a function $f:G\rightarrow\mathbb{R}$.
Figure 10: The evolution of the mean Fourier Entropy per layer over the course of training for 128 runs trained on $S_{5}$. The maximum value $\mathbf{H}[\hat{f}]$ can take over $S_{5}$ is $\log(\frac{1}{7})\approx 1.946$. $\mathbf{H}[\hat{f}]=0$ implies that $\hat{f}(\rho)=0$ on all but one irrep.
In fact, pretty much anywhere in the model where a matrix or set of activations has $|G|$ in the shape we can expand into the Fourier basis. For non-abelian groups, each Fourier frequency is an irrep, and the Fourier transform for each irrep is matrix-valued. This is, on its face, less than interpretable. Following the techniques outlined in Diaconis [12], however, we can expand the function at each element $g\in G$ into a new Fourier basis. Concretely, if our function $f:G\rightarrow\mathbb{R}$ is represented as a vector, we know from 5 that each element of the vector is a sum of the Fourier components:
@@ E.1 Projection on Fourier Basis
We can keep track of all of the Fourier components at once by purposefully not completing the sum from 5), but instead keep each term into a new dimension:
@@
for each coset $xH$. Because Fourier transform $\hat{f}$ is invariant under translation we may, without loss of generality, analyze only the function $f_{H}$. For a given $\alpha_{x}$, $\hat{f}_{xH}(\rho)=\hat{f}^{x}_{H}(\rho)=\rho(x)\hat{f}_{H}(\rho)$ for all $x\in G$. Recall the definition of $\hat{f}_{H}(\rho)$ from E.1:
@@
where in 8 we decompose $\rho|_{H}$ into a direct sum of irreps of $H$. But because each $\tau_{i}$ is irreducible, $\sum_{h\in H}\tau_{i}(h)=\mathbf{0}$ unless $\tau_{i}$ is the trivial irrep. Thus, unless the decomposition of $\rho_{H}$ into irreps of $H$ includes the trivial representation, $\hat{f}|_{H}=0$
E.2 Logits and Counting Cosets
@@
Appendix F The Group Composition via Representations Algorithm
Our work and analysis owes much to ”A Toy Model of Universality,” though we come to different conclusions. Here we go over in detail the four main threads of evidence for the GCR algorithm that Chughtai et al. [6] put forward, which we restate here for clarity: (1) Correlation between the model’s logits and characters of a learned representation $\rho$. (2) The embedding and unembedding layers function as a ”lookup table” for the representations of the input elements $\rho(a),\;\rho(b)$ and the inverse of the target $\rho(c^{-1})$. (3) The neurons in the linear layer calculate the matrix product $\rho(a)\rho(b)=\rho(ab)$. (4) Ablations showing that the circuit they identify is responsible for the majority of the model’s performance. Though we were able to reproduce and verify almost all of Chughtai et al. [6]’s experimental results, we did not find evidence for the GCR algorithm. We instead found evidence that the models were encoding information about coset membership in their weights and, upon careful analysis, realized that the connections between functions from $S_{n}\rightarrow\mathbb{R}$ with low Fourier Entropy and the cosets of $S_{n}$ provides an alternative explanation for the evidence that Chughtai et al. [6] provide.
F.1 Our Interpretation of the Evidence for GCR
Chughtai et al. [6] proposed as a hypothesis the ”Group Composition via Representations” (GCR) algorithm in which the model encodes the group’s irreducible matrix representations (irreps) in its weights and performs the requisite matrix multiplications to solve the task.
Although we replicated almost all of their analytical results, we could not confirm this hypothesis or find further evidence of the GCR algorithm. Close analysis revealed alternative explanations for data such as the observed sparsity in the group Fourier basis and the correlation between the model’s logits and characters of irreps. We show that all of these observations are also consistent with the coset circuit we identify.
Irrep Look Up Table We were not able to find any evidence that the embedding or unembedding layers function as a look-up table for any representation except for the one-dimensional sign representation. We did find that the model’s weights and activations concentrate on specific irreps in the group Fourier basis. This is due, however, to concentration on cosets of specific subgroups, not because the matrix representations are realized anywhere in the weights. The relationship between functions that are constant on cosets and specific irreps is shown in Appendix E.3.
Logit Attribution We also find that the model’s logits correlate with the character $\chi_{\rho}(abc^{-1})$ when the irrep $\rho$ appears in the Fourier transform of the model’s weights. This is not, however, because the model has implemented the matrix product $\rho(ab)\rho(c^{-1})$, but because the model is counting the number of cosets that $ab$ and $c$ are both in. As we prove in E.2, if the cosets are of conjugate subgroups that have their Fourier transform concentrated on the irrep $\rho$ (as we observe for the models in question), then the number of shared cosets will also correlate with the characters of $\rho$.
Matrix Multiplication of Irreps We were not able to find any evidence that the linear layer implements matrix multiplication, again excluding scalar multiplication of the sign irrep. Instead we show empirically that each neuron specializes in recognizing a pair of conjugate subgroups that have a set of shared cosets on which the group multiplication is simpler and predictable.
F.2 Evidence GCR Does Not Explain
The Difference Between Subgroups and Irreps We show that more than one subgroup of $G$ may be have the Fourier transform of its indicator function concentrate on one irrep, such as $S_{n-1}$ concentrating on entirely on $(n-1,1)$. Additionally, as we see for subgroups of $S_{6}$, many subgroups do not concentrate on only one irrep. This makes the assertion that there is a one-to-one relationship between irreps and circuits difficult to support, which at the least complicates GCR’s account of the mechanism.
Quantitatively speaking, the irreps (and therefore the GCR) already fail to catch up with the number of subgroups. Take $S_{n}$ as examples, we have the following table  [^2] The numbers of subgroups use the A005432 sequence of OEIS Foundation Inc. [43]. The numbers of irreps use the A000638 sequence of OEIS Foundation Inc. [43]:
Table 5: The number of subgroups and the number of irreps from $S_{5}$ to $S_{13}$
$S_{5}$ $S_{6}$ $S_{7}$ $S_{8}$ $S_{9}$ $S_{10}$ $S_{11}$ $S_{12}$ $S_{13}$
Number of subgroups $30$ $156$ $1455$ $11300$ $151221$ $1694723$ $29594446$ $404126228$ $10594925360$
Number of irreps $11$ $19$ $56$ $96$ $296$ $554$ $1593$ $3094$ $10723$
Table 5: The number of subgroups and the number of irreps from $S_{5}$ to $S_{13}$
@@ F.2 Evidence GCR Does Not Explain
Asymptotically, the number of subgroups of $S_{n}$ is bounded below as follows (see [48, Corollary 3.3]):
@@
whereas the number of irreps of $S_{n}$ is asymptotically the following (see [17]):
@@
Unembedding Correlations of Neurons We observe that, except in the rare instances when a ”positive” coset neuron forms, the correlation between neurons that concentrate on the same coset is $\approx 100\%$. The correlation between other neurons neurons that concentrate on other cosets but are similar in Fourier space (have correlated irreps) only have $0-50\%$ correlation.
Appendix G A Walkthrough of an $S_{4}$ Circuit
  1. 1.
    In the left embedding layer, encode which right coset of $S^{5}_{4}$ each permutation is in.
  2. 2.
    In the right embedding layer, encode which left coset of $S^{2}_{4}$ each permutation is in.
  3. 3.
    In the left linear layer, query the embeddings for the coset membership of $\sigma_{l}$. The pre-activations must only be dependent on the membership in the right cosets of $S^{5}_{4}$. The left pre-activations are $\sigma_{l}\in S^{5}_{4}\tau_{(i,5)}\mapsto x_{\tau_{i,5}}$.
  4. 4.
    In the right linear layer, query the embeddings for the coset membership of $\sigma_{r}$. The pre-activations must only be dependent on the membership in the left cosets of $S^{2}_{4}$. The right pre-activations are, $\sigma_{r}\in\tau_{(2,j)}S_{4}\mapsto x_{\tau_{(}2,j)}$.
  5. 5.
    The left and right pre-activations are added together, and again there are three options.
    1. (a)
      If $\sigma_{l}\in S_{4}\tau$ and $\sigma_{r}\in\tau S_{4}$ for the same transposition $\tau$, then the pre-activation will be 0.
    2. (b)
      $\sigma_{l}$ and $\sigma_{r}$ are not in matched cosets and $x_{\tau_{l}}+x_{\tau_{r}}>0$. Then the neuron fires and the unembedding layer knows that $\sigma_{l}\sigma_{r}\notin S_{4}$
    3. (c)
      $\sigma_{l}$ and $\sigma_{r}$ are not in matched cosets and $x_{\tau_{l}}+x_{\tau_{r}}<0$. The ReLU will clip the negative pre-activation to 0.
Appendix H Irreducible Representations
H.1 Symmetric Group $S_{5}$

### v2

@@ Acknowledgements
We would like to thank Coreweave for donating the computing resources that we used to run all of our experiments, to Bilal Chughtai for helpful discussions we had throughout the project, and to Neel Nanda for telling us to “not hold back for fear of offending [him].” We would also like to thank Nora Belrose, Neils uit de Bos, Aidan Ewart, Sara Price, Hailey Schoelkopf, Cédric Simal, and Benjamin Wright for their helpful feedback on earlier drafts of this paper.
@@ Appendix A Author Contributions
Wrote the code for training and for calculating the Group Fourier Transform over $S_{n}$. Performed the initial analyses of models trained on $S_{5}$ and initially found what we came to call the coset circuit. Designed and ran causal experiments to confirm our understanding of the coset circuit. Derived formal properties of the coset circuit. Participated in discussions throughout the project and in writing the paper.
@@
Appendix B Structure of the Appendix
In the Appendix we provide more of the mathematical background needed to fully describe some of our results and techniques. In particular, we explain the Group Fourier Transform and how we used to to analyze our models. We do this because we believe it is of independent interest and also because it is necessary to fully explain where our results and those of Chughtai et al. [4] diverge.
In Appendix C we go over the precise experimental set up of the models that we trained.
In Appendix D we introduce the necessary concepts from group theory needed to rigorously talk about the more mathematical aspects of our results.
In Appendix E we introduce representation theory, representations of the symmetric group, and the group Fourier transform.
In Appendix G we return to the coset circuit and coset neurons, with the presentation grounded in the mathematical concepts introduced in Appendices D and E.
Finally, in Appendix H we present extra graphs that did not fit in the main paper and in Appendix I we present a table of all of the conjugacy classes of subgroups of $S_{5}$.
Appendix C Experiment Details
We conducted experiments focusing on the permutation group of $S_{5}$ and $S_{6}$. All models were trained on NVIDIA GeForce RTX 2080 GPUs. All models were implemented in PyTorch Paszke et al. [45] and trained with the Adam optimizer [28] with a fixed learning rate of $0.001$, weight decay set to $1.0$, $\beta_{1}=0.9$ and $\beta_{2}=0.98$. At the beginning of each training run, the training set is sampled uniformly from all $|S_{n}|^{2}$ combinations of permutations. Each optimization step was made on the entire training set. Using our setup a single $S_{5}$ model trained in approximately 8 hours and a single $S_{6}$ model trained in approximately 100 hours, though multiple training jobs could be scheduled on a single GPU. Analysis and reverse engineering was performed with Vink et al. [56], Nanda & Bloom [39], Harris et al. [22], GAP [16], Stein et al. [53].
Table 2: Experiment hyperparameters.
Group % Train Set Num. Runs Num. Epochs Linear Layer Size Embedding Size
$S_{5}$ 40% 128 250,000 128 256
$S_{6}$ 40% 100 50,000 256 512
@@ D.3 Normal Subgroups
The double cosets of a normal subgroup are simply the usual cosets.
@@ D.5 An important case
Appendix E Representation Theory
E.1 Preliminaries
Definition E.1.
@@ E.1 Preliminaries
In other words, a representation maps a group element $g$ to a linear operator $f(g):V\rightarrow V$ where $V$ is a vector space of dimension $d$, so that the group multiplication becomes compositions of linear operators ($f(g\cdot h)=f(g)\circ f(h)$). Without explicit specifications, all representations in this paper are assumed to be over complex numbers. Recall also that finite dimensional linear operators can be represented as matrices, and composition of linear operators is then given as matrix multiplication.
@@
Example E.10.
The permutation representation is reducible, because there is a subspace of $\mathbb{C}^{3}$ that is invariant to it’s action.
@@
As it turns out, there are no irreducible representations of $S_{3}$ that are three-dimensional. The largest irrep of $S_{3}$ is $\rho_{(2,1)}$, which is made of $2\!\times\!2$ matrices. The matrices of the $(2,1)$ irrep of $S_{3}$ are as follows:
(1 2 3)123(1\;2\;3)( 1 2 3 ) $\mapsto\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$ (2 1 3)213(2\;1\;3)( 2 1 3 ) $\mapsto\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}$ (3 2 1)321(3\;2\;1)( 3 2 1 ) $\mapsto\begin{pmatrix}1/2&-\sqrt{3}/2\\ \sqrt{3}/2&-1/2\end{pmatrix}$
(1 3 2)132(1\;3\;2)( 1 3 2 ) $\mapsto\begin{pmatrix}-1/2&\sqrt{3}/2\\ \sqrt{3}/2&1/2\end{pmatrix}$ (3 1 2)312(3\;1\;2)( 3 1 2 ) $\mapsto\begin{pmatrix}-1/2&\sqrt{3}/2\\ -\sqrt{3}/2&-1/2\end{pmatrix}$ (2 3 1)231(2\;3\;1)( 2 3 1 ) $\mapsto\begin{pmatrix}-1/2&\sqrt{3}/2\\ -\sqrt{3}/2&-1/2\end{pmatrix}$
@@
Appendix F Fourier transform over finite groups
@@
For more details and applications, one can refer to, for example, Elias M. Stein [11]. We would like to note that there is also an inverse transform that restores the original function $f$ from $\hat{f}$:
@@
Appendix G The Coset Circuit (with more math)
We did not introduce it in the main body of our paper because it would distract from the core of our results, but for the first half of our investigation the Fourier transform over the symmetric group was integral to our investigation. We were building directly on [4] who had shown striking results around the weights of single-layer models showing high degrees of correlation with the irreps of the symmetric group. We wished to cast those results in the language of the group Fourier transform. Even when we realized that the mechanism of the model was based around cosets it became extremely important to understand why our coset circuit was so concentrated in Fourier space.
G.1 Harmonic Analysis on the Symmetric Group
The presentation in the Appendix E was given in terms of functions on $\mathbb{C}$ because it is required for arbitrary groups. For $S_{n}$ all of the irreps are rational [15] and the Fourier transform of functions on $S_{n}$ can safely be defined over $\mathbb{R}$.
In this section we describe how we use the Fourier transform to analyze the weights and activations of an MLP. The inputs to the model are two one-hot vectors, $\mathbf{x}_{l},\;\mathbf{x}_{r}$, which multiply the embedding matrices $\mathbf{E}_{l}\mathbf{x}_{l}$ and $\mathbf{E}_{r}\mathbf{x}_{r}$. $\mathbf{E}_{l}$ and $\mathbf{E}_{r}$ are $d\times|G|$ matrices, where $d$ is the embedding dimension and $|G|$ is the size of the group. The columns are the embedding vectors for a single element $g\in G$. The normal approach would be to try and look at the column spaces of $\mathbf{E}_{l}$ and $\mathbf{E}_{r}$, as these columns are the inputs to the model. However, since each row of $\mathbf{E}_{l}$ and $\mathbf{E}_{r}$ and each value of that row is associated with a single element of $G$, we instead treat each row of the embedding as a function $f:G\rightarrow\mathbb{R}$.
In fact, anywhere in the model where a matrix or set of activations has $|G|$ in the shape we can expand into the Fourier basis. For non-abelian groups, each Fourier frequency is an irrep, and the Fourier transform for each irrep is matrix-valued. This is, on its face, less interpretable than what we started with. Following the techniques outlined in Diaconis [8], however, we can expand the function at each element $g\in G$ into a new Fourier basis. Concretely, if our function $f:G\rightarrow\mathbb{R}$ is represented as a vector, we know from 6 that each element of the vector is a sum of the Fourier components:
@@ G.1 Harmonic Analysis on the Symmetric Group
We can keep track of all of the Fourier components at once by purposefully not completing the sum from 6), but instead keep each term into a new dimension:
@@
for each coset $xH$. Because Fourier transform $\hat{f}$ is invariant under translation we may, without loss of generality, analyze only the function $f_{H}$. For a given $\alpha_{x}$, $\hat{f}_{xH}(\rho)=\hat{f}^{x}_{H}(\rho)=\rho(x)\hat{f}_{H}(\rho)$ for all $x\in G$. Recall the definition of $\hat{f}_{H}(\rho)$ from F.1:
@@
where in 9 we decompose $\rho|_{H}$ into a direct sum of irreps of $H$. But because each $\tau_{i}$ is irreducible, $\sum_{h\in H}\tau_{i}(h)=\mathbf{0}$ unless $\tau_{i}$ is the trivial irrep. Thus, unless the decomposition of $\rho_{H}$ into irreps of $H$ includes the trivial representation, $\hat{f}|_{H}=0$
G.2 Logits and Counting Cosets
@@
Indeed, we see that when looking at the action of an individual neuron on the prediction space (i.e. “if this neuron fires, which predictions become more likely and which less?”), we see that it is only neurons that are predicting the same coset that are correlated. The average pairwise correlation of neuron actions is uncorrelated, as is the correlation of neurons associated with the same irrep. Refer to Table 3 for the full results.
Table 3: The correlation of unembedding neurons. Neurons that correspond to the same coset are averaged together in the unembedding, leading to the unembedding vectors being highly correlated.
Mean Correlation Std Dev Correlation
Within Coset 0.814 0.445
Within Subgroup Conjugacy Class -0.002 0.222
Baseline -0.003 0.163
G.3 An Asymptotic Analysis
Our theory of coset circuits and the GCR algorithm of [4] cannot be equivalent because there is no one-to-one relationship between irreps and subgroups. Even for $S_{5}$, there are more subgroups than irreps. Quantitatively speaking, the irreps already fail to catch up with the number of subgroups. For the direct comparison of $S_{n}$ refer to
@@ G.3 An Asymptotic Analysis
Asymptotically, the number of subgroups of $S_{n}$ is bounded below as follows (see Pyber [48, Corollary 3.3]):
@@
whereas the number of irreps of $S_{n}$ is asymptotically the following (see Erdos [13]):
@@
In practice, as can be seen in Table 5, many subgroups concentrate on more than one irrep. We do not have an explanation for why the coset circuits always do concentrate one irrep. In practice, the different values for the cosets are arranged so that the contributions of all but one irrep cancel out. We hypothesize that it may have something to do with the margin maximization effect discussed in [38]. As we mention in the main body, we observe that subgroups which concentrate on more than one irrep will form coset circuits that concentrate entirely on any of the irreps, while still behaving equivalently. We do not think that there is in fact a connection between what the circuit is doing the irrep.
Table 4: The number of subgroups and the number of irreps from $S_{5}$ to $S_{12}$. The numbers of subgroups use the A005432 sequence of the OEIS [42]. The numbers of irreps corresponds to the number of integer partitions of $n$ and use the A000041 sequence of the OEIS [42].
$S_{5}$ $S_{6}$ $S_{7}$ $S_{8}$ $S_{9}$ $S_{10}$ $S_{11}$ $S_{12}$
Number of subgroups $156$ $1455$ $11300$ $151221$ $1694723$ $29594446$ $404126228$ $10594925360$
Number of irreps $7$ $11$ $15$ $22$ $30$ $42$ $56$ $77$
Appendix H Extra Graphs
H.1 Distribution over Subgroups and Cosets
(a) 128 Models trained on $S_{5}$
(b) 100 Models trained on $S_{6}$
Figure 6: Distribution of coset circuits for models trained on $S_{5}$ and $S_{6}$ with different initial seeds. Every model has a few sign circuit neurons that correspond to $A_{n}<S_{n}$, but the model cannot completely solve the task with only the sign circuit, so there are never more than a few. Every other subgroup could, with enough neurons, be used to completely solve the the multiplication, but in general if a model primarily uses a single subgroup it is $S_{n-1}$ (in the main body of the paper we refer to these subgroups as $H_{i}$, for the element $i\in[n]$ that is fixed). Every model has at least a few $S_{n-1}$ neurons. Many models use a mix of subgroups and there is often a “long tail” of a subgroup being represented by only one or two neurons. The subgroups marked with asterisks, $A_{5}^{*}$ and $S_{5}^{*}$, correspond to the “exceptional” subgroups of $S_{6}$, which come from an outer automorphism that only $S_{6}$ has [26]. These subgroups are isomorphic to $S_{5}$ and $A_{5}$, but not conjugate to the subgroups that come from fixing an element in $\{1,\dots,6\}.$
H.2 Other Examples of Coset Circuits Forming
(a) $S_{3}\!\times\!S_{2}$ Left Permutations
(b) $S_{3}\!\times\!S_{2}$ Right Permutations
Figure 7: The formation of an $S_{3}\!\times\!S_{2}$ neuron.
(a) $A_{4}$ Left Permutations
(b) $A_{4}$ Right Permutations
Figure 8: The formation of an $A_{4}$ neuron.
Appendix I Irreducible Representations
I.1 Symmetric Group $S_{5}$